OpenStudy (anonymous):

When solving a rational equation what is the first step we must always take?

hero (hero):

Find LCD or get rid of denominators

OpenStudy (anonymous):

rationalize the denominator

OpenStudy (anonymous):

...by multiplying the equation by "1", the fraction of the rational divided by the rational.

OpenStudy (anonymous):

would i square anything?

OpenStudy (anonymous):

Get rid of the denominator.

OpenStudy (anonymous):

Not really. For instance, in \[4/\sqrt{3}\], you would rationalize it by multiplying th fraction by \[\sqrt{3}/\sqrt{3}\], so it's similar to squaring the denominator. You would multiply the numerators, however, getting 4√3 in the numerator. In the denominator √3 times √3 is 3.

OpenStudy (anonymous):

depends on the equation, for example √25=√(x+1) you would square it in this case 25=x+1 x=24